Zari:

3,9,7,13,11,17,15 what are the next four pattern below ?

Jaynater:

Try to look at the difference between numbers from one to the next, and what do you notice?

Zari:

21,23,17,19 13,11,17,23 13,19,17.23 21,19,25,23

umm:

If you look at the pattern, it continues as +6 -2 +4 and again. Which one do you think it may be?

Zari:

@umm wrote:
If you look at the pattern, it continues as +6 -2 +4 and again. Which one do you think it may be? tbh I think its 13,19,17,23

Icyhot:

@zari wrote:
@icyhot
I suck at Maths

Zari:

@icyhot wrote:
@zari wrote:
@icyhot
I suck at Maths oh lord

snowflake0531:

+6 -2 3,9,7,13,11,17,15 3+6=9 9=2=7 etc 17-2=15 so then you should add 6 to 15

Zari:

@snowflake0531 wrote:
+6 -2 3,9,7,13,11,17,15 3+6=9 9=2=7 etc 17-2=15 so then you should add 6 to 15 okay

snowflake0531:

so what number comes after 15

Zari:

@snowflake0531 wrote:
so what number comes after 15 20

snowflake0531:

stop quoting me each time and no, 15+6 is not 20

Zari:

@snowflake0531 wrote:
stop quoting me each time and no, 15+6 is not 20 I meant 21*

snowflake0531:

okay, so you can eliminate two answers now, after adding 6, you need to subtract two what is 21-2

Zari:

19

snowflake0531:

so the following pattern would be 21 then 19 you have your answer now

Zari:

okay